# Digital Twin Calibration for Biological System-of-Systems: Cell Culture Manufacturing Process

### v1

@@ Abstract
Biomanufacturing innovation relies on an efficient design of experiments (DoE) to optimize processes and product quality. Traditional DoE methods, ignoring the underlying bioprocessing mechanisms, often suffer from a lack of interpretability and sample efficiency. This limitation motivates us to create a new optimal learning approach that can guide a sequential DoEs for digital twin model calibration. In this study, we consider a multi-scale mechanistic model for cell culture process, also known as Biological Systems-of-Systems (Bio-SoS), as our digital twin. This model with modular design, composed of sub-models, allows us to integrate data across various production processes. To calibrate the Bio-SoS digital twin, we evaluate the mean squared error of model prediction and develop a computational approach to quantify the impact of parameter estimation error of individual sub-models on the prediction accuracy of digital twin, which can guide sample-efficient and interpretable DoEs.
@@ 1 INTRODUCTION
To support interpretable predictions and optimal control of biomanfuacturing processes, in this paper, we develop a digital twin calibration approach to guide a sequential design of experiments (DoEs). It can sample efficiently improve model fidelity of multi-scale bioprocess mechanistic model for Biological System-of-Systems (Bio-SoS) (Zheng et al., 2024) that characterizes the fundamental bioprocessing mechanisms and causal interdependence from molecular- to cellular- to macro-kinetics. Even though this study is motivated by cell culture process, the proposed optimal learning approach can be extended to calibrate the digital twin of general Bio-SoS with modular design. In specific, the dynamics and variations of cell culture process is characterized by a multi-scale mechanistic model composed of sub-modules or modules: (1) a single cell mechanistic model characterizing each living cell behaviors and their interactions with environment; (2) a metabolic shift model characterizing the change of cell metabolic phase and behaviors as a response to culture condition changes and cell age; and (3) macro-kinetic model of a bioreactor system composed of many living cells under different metabolic phases.
@@
The benefits of considering the Bio-SoS mechanistic model with modular design include: a) support flexible manufacturing through assembling a system of modules to account for biomanufacturing processes under different conditions and with various inputs; and b) facilitate the integration of experimental data from different production systems, such as 2D culture and 3D culture for Induced pluripotent stem cell (iPSCs). By accounting for the structure property of the Bio-SoS mechanistic model, we can quantify how the model uncertainties or estimation errors of different modules interact with each other and propagate through the reaction mechanism pathways to outputs, which can support the interpretable design of experiments to enhance model fidelity and improve predictions.
@@
Before describing the proposed calibration approach, we review the key aspects for digital twin calibration. First, the model uncertainty quantification approaches for digital twins and simulators can be divided into two main categories: Bayesian and frequentist approaches (Corlu et al., 2020). Bayesian approaches treat unknown model parameters as random variables and quantify our belief by using posterior distributions. It involves specifying prior distributions for model parameters and updating these distributions based on the information collected from observed data by applying Bayes’ theorem. On the other hand, frequentist approaches estimate model parameters using traditional statistical methods, such as maximum likelihood estimation, to find parameter values that maximize the likelihood of observed data. The model estimation uncertainty can be quantified by asymptotic normal approximation or bootstrap approach. These frequentist estimation approaches could be easier and faster compared with Bayesian inference approaches since there may not exist conjugate priors and posterior sampling update can be computationally expensive.
@@
Second, for the propagation of input and model uncertainty to simulation outputs, to save computational budget and time, a metamodel is often used to approximate the response surface, especially for complex stochastic systems, since easy experiment can be expensive. In the classical calibration approach (Kennedy and O’Hagan, 2001), the Gaussian processes (GPs) are used to characterize the mean response and model discrepancies. Although GP metamodel is often used to propagate the input to output and further support the digital twin calibration, it’s hard to interpret and difficult to leverage prior knowledge of the real biomanufacturing system mechanisms. To enhance interpretability and sample efficiency, the mechanistic models can be employed to construct the response surface in our study.
@@
Third, the selection of calibration criteria plays a critical role to guide the experiments and most informative data collection to improve model fidelity and support digital twin calibration. The existing calibration approaches often focus on system design. Mean response and Mean Squared Error (MSE) are widely used as a criteria for assessing model prediction accuracy. Tuo and Jeff Wu (2016) shows Kennedy’s method, which models the mean output as a GP, may lead to asymptotically $L_{2}$-inconsistent. They modify the criteria into $L_{2}$ norm of the discrepancy between real and digital system outputs and propose a $L_{2}$-consistent calibration method, which can ensure an optimal convergence rate.
@@
To support process control and facilitate biomanufacturing process automation, in this paper, we aim to calibrate the Bio-SoS mechanistic model to guide a sequential DoE and improve model fidelity and prediction accuracy of digital twin. We first employ the Maximum Likelihood Estimation (MLE) method to estimate model parameters and utilize bootstrap techniques to quantify estimation uncertainty across different modules. Then, to optimize the digital twin calibration policy that will be used to sequentially guide the experiments and the most informative data collection, we derive the gradient of prediction MSE and use the steepest descent to efficiently search for the optimal policy parameters, which makes the learning process more interpretable. Since it is challenging to find a closed form solution of the mechanistic model with stochastic differential equations (SDEs) form, we utilize the linear noise approximation (LNA) for uncertainty propagation and constructs a surrogate model for MSE. Thus, in an online setting, a calibration policy is iteratively updated by using a stochastic gradient method where the gradient of MSE with respect to the calibration policy parameter follows the backward direction of the uncertainty propagation. The proposed calibration approach accounts for the Bio-SoS mechanism structure. That means it quantifies Bio-SoS module error interaction and their propagation through mechanistic pathways to output, which guides the sequential design of experiments to efficiently improve model fidelity and prediction accuracy.
@@
In sum, we develop an interpretable and sample efficient optimal learning approach to support digital twin calibration for a multi-scale bio-process mechanistic model so that the digital twin of the cell culture process can improve process prediction and support optimal control. Even though this paper focuses on cell culture, the proposed calibration approach can be extendable to general biomanufacturing systems. The key contributions of the proposed calibration approach and the benefits are summarized as follows.
@@
  • We proposed a new DoE method for calibrating a multi-scale bioprocess mechanistic model by updating the parameters using closed-form gradients that incorporate the mechanistic information.
@@
  • We assess the MSE of the process output prediction and use a LNA-based metamodel along with Euler’s method to estimate how model uncertainty propagates through Bio-SoS mechanism pathways and impacts on the output prediction accuracy. This approach can advance our understanding on how the errors in individual module parameters affect the overall prediction accuracy of digital twin model.
@@
  • Built on the LNA, we further develop a gradient-based policy optimization to guide most informative experiments and support sample-efficient optimal learning.
@@
The organization of the paper is as follows. In Section 2, we describe the multi-scale mechanistic model. Then, we propose the novel calibration method for the multi-scale model in Sections 3, which includes two main procedures: model inference and policy update. A stochastic gradient method is developed for the model inference in Section 4. The calibration policy gradient estimation and update method is further developed in Section 5. Then we empirically validate the proposed digital twin calibration approach in Section 6 and conclude this paper in Section 7.
@@ 2 Problem Description and Bio-SoS Mechanistic Model
In this study, we explore a multi-scale bioprocess mechanistic model designed to elucidate the interdependencies that exist across molecular, cellular, and macroscopic levels within cell culture processes. This comprehensive model integrates the complexities of biological interactions within and between these scales and extends applicably to a broader range of biological systems, referred to as Biological System-of-Systems (Bio-SoS). A detailed summary of the variable notations used within our model is provided in Table 1.
The core of our model is based on the structured Markov chain, capturing both the dynamics and variability inherent in cellular processes through a state transition probabilistic model. Conceptually, each individual cell operates as a complex system, and collectively, numerous cells at various metabolic phases within the bioreactor form an intricate system of systems. This interaction is depicted in Figure 1, where, for clarity, only three distinct metabolic phases are illustrated. Cells interact with each other by altering their environment, notably through nutrient uptake and the production of metabolic wastes, and in turn, respond to these environmental changes.
@@
At any time $t$, for each single cell in the $i$-th metabolic phase, denoted by $Z_{t}=i$ with $i\in\{0,1,\ldots,I\}$ (i.e., growth, stationary, death phases, etc), the dynamic behaviors of its metabolic network can be characterized by a stochastic model specified by parameters $\boldsymbol{\alpha}_{i}$. The regulation mechanism and reaction rate of metabolic network respond to the change of environmental condition is denoted by $\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. The metabolic phase shift model with $\mathbb{P}_{ij}(\boldsymbol{s}_{t};\boldsymbol{\beta})$ represents the transition probability from $Z_{t}=i$ to $Z_{t}=j$. Therefore, in this paper, we focus on calibrating the mechanistic model of the Bio-SoS specified by the calibration parameters $\boldsymbol{\theta}\equiv\{\boldsymbol{\alpha}_{i},\boldsymbol{\beta}\}$: (1) $\boldsymbol{\alpha}_{i}$ characterizing cell dynamics in $i$-th metabolic phase or class; and (2) $\boldsymbol{\beta}$ characterizing the phase shift probability that impacts the percentage of cells in each phase or class.
@@ 2.1 Single Cell Model
This section briefly describes a stochastic metabolic model characterizing the dynamics of a single cell across different metabolic phases: growth ($Z=0$), production ($Z=1$) and so on. The model includes two main components: the Stochastic Modular Reaction Networks (SMRN) and the metabolic phase shift dynamics. Following the study (Anderson and Kurtz, 2011), the SMRN can be constructed
$\tilde{\mathbf{s}}^{i}_{t+1}=\tilde{\mathbf{s}}^{i}_{t}+\mathbf{N}\cdot\boldsymbol{R}_{s}^{i}(t),$ (1)
where $\tilde{\mathbf{s}}^{i}_{t}$ is the state for cells in $i$-th phase, $\mathbf{N}$ represents the stoichiometric matrix characterizing the structure of molecular reaction network, and $\boldsymbol{R}_{s}^{i}(t)$ follows a multivariate Poisson process with molecular reaction rate vector $\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. This reaction rate depends on macro-environment $\boldsymbol{s}_{t}$, typically modeled by using a Michaelis-Menten (MM) kinetic equation (Kyriakopoulos et al., 2018; Wang et al., 2024).
The second component on metabolic shift model addresses the transitions between metabolic phases in response to environmental changes. At any time $t$, we have the metabolic phase-shift probability matrix $\mathbb{P}(\boldsymbol{s}_{t};\boldsymbol{\beta})$ depending on the environmental condition $\boldsymbol{s}_{t}$, where each element represents
@@ 2.2 Macro-Kinetic State Transition
where $\mu_{i}{(\boldsymbol{s}_{t})}$ represents growth rate characterizing the growth of the cell population and the second term counts the cells metabolic shifting from $i^{\prime}$-th phase to $i$-th phase.
Let $\boldsymbol{s}_{t}$ denote metabolite concentrations (i.e., the number of molecules per unit of volume) in the system at time $t$. Let $\Delta\tilde{\mathbf{s}}_{t+1}^{(i,n)}=\tilde{\mathbf{s}}^{(i,n)}_{t+1}-\tilde{\mathbf{s}}^{(i,n)}_{t}$ represent the change in metabolic concentration for the $n$-th cell in phase $i$. Then at any time $t$, given the density $X_{i,t}$ of cells located in phase $i$, the overall change of metabolite concentration is the sum of the contributions from individual cells,
@@ 2.3 State Transition Probability
In an infinitesimal time change $\mathrm{d}t$, the change of state $\boldsymbol{s}_{t}$ becomes,
$\mathrm{d}\boldsymbol{s}_{t}=\boldsymbol{s}_{t+\text{d}t}-\boldsymbol{s}_{t}=\sum_{i=0}^{1}\mathbf{N}\text{d}\boldsymbol{R}^{i}(t)=\mathbf{N}\text{d}\boldsymbol{R}(t),$
where $\boldsymbol{R}(t)$ is a multivariate Poisson process with parameter $\sum_{i=0}^{1}X_{i,t}\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. Then ${\mathbb{E}}(\text{d}\boldsymbol{R}\mid\boldsymbol{s}_{t})=\sum_{i=0}^{1}X_{i,t}\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$ and ${\rm Var}(\text{d}\boldsymbol{R}\mid\boldsymbol{s}_{t})\approx\text{diag}\{\sum_{i=0}^{1}X_{i,t}\boldsymbol{\nu}^{r}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})\}$, with $r=1,2,\cdots,R$ representing the component indices of $\boldsymbol{\nu}$. The approximation is made under the assumption that the time interval is small, so the state doesn’t change too much and different reactions can be treated independent. By forming as a Stochastic Differential Equations (SDEs), we have
@@
where $\boldsymbol{B}_{t}$ denotes standard Brownian motion. It is hard to find the analytic solution to (4), so we take an approximation in a short time interval $(t,t+\Delta t]$, i.e.,

### v2

@@ Abstract
Biomanufacturing innovation relies on an efficient Design of Experiments (DoEs) to optimize processes and product quality. Traditional DoE methods, ignoring the underlying bioprocessing mechanisms, often suffer from a lack of interpretability and sample efficiency. This limitation motivates us to create a new optimal learning approach for digital twin model calibration. In this study, we consider the cell culture process multi-scale mechanistic model, also known as Biological System-of-Systems (Bio-SoS). This model with a modular design, composed of sub-models, allows us to integrate data across various production processes. To calibrate the Bio-SoS digital twin, we evaluate the mean squared error of model prediction and develop a computational approach to quantify the impact of parameter estimation error of individual sub-models on the prediction accuracy of digital twin, which can guide sample-efficient and interpretable DoEs.
@@ 1 INTRODUCTION
To support interpretable predictions and optimal control of biomanfuacturing processes, in this paper, we develop a digital twin calibration approach for multi-scale bioprocess mechanistic model or Biological System-of-Systems (Bio-SoS) (Zheng et al., 2024) characterizing causal interdependence from molecular- to cellular- to macro-kinetics. Even though this study is motivated by cell culture process, it can be extended to calibrate general Bio-SoS with modular design. Basically, cell culture process dynamics and variations depend on the modules: (1) a single cell mechanistic model characterizing each living cell behaviors and their interactions with environment; (2) a metabolic shift model characterizing the change of cell metabolic phase and behaviors as a response to culture conditions and cell age; and (3) macro-kinetic model of a bioreactor system composed of many living cells under different metabolic phases.
@@
The benefits of considering the Bio-SoS mechanistic model with modular design include: a) support flexible manufacturing through assembling a system of modules to account for biomanufacturing processes under different conditions and inputs; and b) facilitate the integration of heterogeneous data from different production processes, such as 2D culture and 3D aggregate culture for Induced Pluripotent Stem Cells (iPSCs) (Wang et al., 2024; Zheng et al., 2024). By incorporating the structure property of the Bio-SoS mechanistic model into the calibration method, we can quantify how the model uncertainties or approximation errors of different modules interact with each other and propagate through the reaction pathways to the prediction of outputs (e.g., yield and product quality attributes), which can guide interpretable and most informative Design of Experiments (DoEs) to efficiently improve model fidelity with less experiments.
@@
The model uncertainty quantification approaches for digital twin calibration can be divided into two main categories: Bayesian and frequentist approaches (Corlu et al., 2020). Bayesian approaches treat unknown model parameters as random variables and quantify our belief by posterior distributions. It involves specifying prior distributions for model parameters and updating these distributions based on the information from observed data by applying Bayes’ theorem. On the other hand, frequentist approaches rely on traditional statistical methods, such as maximum likelihood estimation, to find parameter values that maximize the likelihood of observed data. Those estimation approaches could be faster under the situations when there not exist conjugate priors and posterior sampling update is computationally expensive.
@@
For the propagation of input and model uncertainty to simulation outputs, to save computational budget and time, a metamodel is often used to approximate the response surface, especially for complex stochastic systems. In the classical calibration approach (Kennedy and O’Hagan, 2001), Gaussian Processes (GPs) are often used to model the mean response and discrepancies, deriving the posterior distribution for predictions and guiding the design of experiments. Although GP metamodel is often used to propagate from inputs to outputs (e.g., mean response), it’s hard to interpret and difficult to leverage prior knowledge of the real system mechanisms (such as bioprocessing reaction network structure). To enhance interpretability and sample efficiency, the mechanistic model structure will be employed to construct the response surface in our study.
@@
Calibration criteria play a critical role to guide DoEs for digital twin calibration. Mean response and Mean Squared Error (MSE) are widely used criteria for assessing model prediction accuracy. Tuo and Jeff Wu (2016) shows Kennedy’s method, which models the mean output as a Gaussian process, may lead to asymptotically $L_{2}$-inconsistent. They modify the criteria into $L_{2}$ norm of the discrepancy between two system outputs and propose a $L_{2}$-consistent calibration method, which has an optimal convergence rate.
@@
To support process prediction and optimal control, in this paper, we aim to calibrate the Bio-SoS mechanistic model with a modular design to improve its predictive accuracy and reduce the MSE of the process output prediction. We employ the Maximum Likelihood Estimation (MLE) method to estimate model parameters and utilize bootstrap techniques to quantify estimation errors across different modules. To optimize the digital twin calibration policy and guide the most informative data collection, we derive a gradient-based approach that follows the steepest descent search in policy parameter space, making the learning process more interpretable. This strategy utilizes the Linear Noise Approximation (LNA) for uncertainty propagation and constructs a surrogate model for MSE. Concurrently, we preserve mechanistic information by solving the Ordinary Differential Equations (ODEs) from LNA using Euler’s method. In an online setting, a calibration policy is iteratively updated by using a stochastic gradient method where the gradient of MSE with respect to the calibration policy parameter follows the backward direction of the uncertainty propagation. The proposed calibration approach accounts for the Bio-SoS mechanism structure and quantifies Bio-SoS module error interaction and their propagation through mechanistic pathways to output prediction, which guides the optimal sequential DoEs to efficiently improve model fidelity and prediction accuracy.
@@
In sum, we develop an interpretable and sample efficient calibration approach for a multi-scale bio-process mechanistic model so that the digital twin of the cell culture process can improve process prediction and support optimal control. Even though this paper uses cell culture as a motivation example of Bio-SoS, the proposed calibration approach can be extendable to general biomanufacturing systems. The key contributions of the proposed calibration approach and the benefits are summarized as follows.
@@
  • We proposed a new sequential DoE method for calibrating a multi-scale bioprocess mechanistic model by updating the parameters using closed-form gradients that incorporate the mechanistic information.
@@
  • We assess the MSE of the process output prediction and use a LNA-based metamodel along with Euler’s method to estimate how model uncertainty propagates through Bio-SoS mechanism pathways and impacts on the output prediction accuracy. This approach can advance our understanding on how the errors in individual module parameters affect the overall accuracy of digital twin model prediction.
@@
  • Built on the LNA, we further develop a gradient-based policy optimization to guide most informative dsign of experiments and support sample-efficient optimal learning.
@@
The organization of the paper is as follows. In Section 2, we describe the multi-scale mechanistic model. Then, we propose the novel calibration method for the multi-scale model in Sections 3, which includes two main procedures: model inference and policy update. A stochastic gradient method is developed for the model inference in Section 4. The calibration policy gradient estimation and update method is further developed in Section 5. Then we empirically validate our method in Section 6 and conclude this paper in Section 7.
@@ 2 Problem Description and Bio-SoS Mechanistic Model
In this study, we explore a multi-scale bioprocess mechanistic model designed to elucidate the interdependencies that exist across molecular, cellular, and macroscopic levels within a cell culture process. This comprehensive model integrates the complexities of biological interactions within and between these scales and extends applicably to a broader range of biological systems, referred to as biological system-of-systems (Bio-SoS). A detailed summary of the variable notations used within our model is provided in Table 1.
Our model is based on the structured Markov chain, capturing both the dynamics and variability inherent in cellular processes through a state transition probabilistic model. Each individual cell operates as a complex system, and collectively, numerous cells at various metabolic phases within the bioreactor form an intricate system of systems. This interaction is depicted in Figure 1, where three distinct metabolic phases are considered. Cells interact with each other by altering their environment, through nutrient uptake and the production of metabolic wastes, and in turn, respond to these environmental changes.
@@
At any time $t$, for each single cell in the $i$-th metabolic phase, denoted by $Z_{t}=i$ with $i\in\{0,1,\ldots,I\}$ (i.e., growth, stationary, death phases, etc), the dynamic behaviors of its metabolic network can be characterized by a stochastic model specified by parameters $\boldsymbol{\alpha}_{i}$ and a metabolic phase shift model with $\mathbb{P}_{ij}(\boldsymbol{s}_{t};\boldsymbol{\beta})$ representing the transition probability from $Z_{t}=i$ to $Z_{t}=j$. Therefore, in this paper, we focus on calibrating the mechanistic model of the Bio-SoS specified by the calibration parameters $\boldsymbol{\theta}\equiv\{\boldsymbol{\alpha}_{i},\boldsymbol{\beta}\}$: (1) $\boldsymbol{\alpha}_{i}$ characterizing cell dynamics in $i$-th metabolic phase or class; and (2) $\boldsymbol{\beta}$ characterizing the phase shift probability that impacts the percentage of cells in each phase or class.
@@ 2.1 Single Cell Model
This section describes a stochastic metabolic model to understand the dynamics of a single cell across different metabolic phases: growth ($Z=0$), production ($Z=1$) and so on. The model includes two main components: the Stochastic Modular Reaction Networks (SMRN) and the metabolic phase shift dynamics. Following the study Anderson and Kurtz (2011), the SMRN can be constructed $\tilde{\mathbf{s}}^{i}_{t+1}=\tilde{\mathbf{s}}^{i}_{t}+\mathbf{N}\cdot\boldsymbol{R}_{s}^{i}(t)$, where $\tilde{\mathbf{s}}^{i}_{t}$ is the state for cells in $i$-th phase, $\mathbf{N}$ represents the stoichiometric matrix characterizing the structure of molecular reaction network, $\boldsymbol{R}_{s}^{i}(t)$ follows a multivariate Poisson process with molecular reaction rate vector $\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. This reaction rate depends on macro-environment $\boldsymbol{s}_{t}$, typically modeled using a Michaelis-Menten (MM) equation (Kyriakopoulos et al., 2018; Wang et al., 2024).
The metabolic shift model addresses the transitions between metabolic phases based on environmental changes. At any time $t$, we have the metabolic phase-shift probability matrix $\mathbb{P}(\boldsymbol{s}_{t};\boldsymbol{\beta})$ depending on the environmental condition $\boldsymbol{s}_{t}$, where each element represents
@@ 2.2 Macro-Kinetic State Transition
where $\mu_{i}{(\boldsymbol{s}_{t})}$ represents cell growth rate and the second term represents the instantaneous cells metabolic shifting from $i^{\prime}$-th phase to $i$-th phase. Let $\boldsymbol{s}_{t}$ denote metabolite concentrations (i.e., the number of molecules per unit of volume) in the system at time $t$. Define $\Delta\tilde{\mathbf{s}}_{t+1}^{(i,n)}$ as the change in metabolic concentration for the $n$-th cell in phase $i$, calculated as $\tilde{\mathbf{s}}^{(i,n)}_{t+1}-\tilde{\mathbf{s}}^{(i,n)}_{t}$. Then at any time $t$, given the density $X_{i,t}$ of cells in phase $i$, the overall change of metabolite concentration is the sum of the contributions from individual cells,
@@ 2.3 State Transition Probability
In a short time interval $[t,t+\mathrm{d}t]$, the change of state $\boldsymbol{s}_{t}$ is denoted by $\mathrm{d}\boldsymbol{s}_{t}=\boldsymbol{s}_{t+\text{d}t}-\boldsymbol{s}_{t}=\sum_{i=0}^{1}\mathbf{N}\text{d}\boldsymbol{R}^{i}(t)=\mathbf{N}\text{d}\boldsymbol{R}(t),$ where $\boldsymbol{R}(t)$ is a multivariate Poisson process with parameter $\sum_{i=0}^{1}X_{i,t}\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. Then ${\mathbb{E}}(\text{d}\boldsymbol{R}\mid\boldsymbol{s}_{t})=\sum_{i=0}^{1}X_{i,t}\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$ and ${\rm Var}(\text{d}\boldsymbol{R}\mid\boldsymbol{s}_{t})\approx\text{diag}\{\sum_{i=0}^{1}X_{i,t}\boldsymbol{\nu}^{r}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})\}$, with $r=1,2,\cdots,R$ representing the component indices of $\boldsymbol{\nu}$. The approximation is made under the assumption that the time interval is small, so the state doesn’t change too much and different reactions can be treated independent. By forming as a Stochastic Differential Equations (SDEs), we have
@@
where $\boldsymbol{B}_{t}$ denotes standard Brownian motion. By Euler-Maruyama method (Bayram et al., 2018), a normal approximation can be applied to (3)